OpenStudy (anonymous):

find the the term in the geometric sequence nth term if a6=27, r=3

OpenStudy (dan815):

a1,a2,a3,a4,a5,A6=27,a7,a8... sooo 27 is the 6th term, we have to write an equation for any term

OpenStudy (dan815):

6th term we got by saying a6=a1*r^5 so An an=a1*r^(n-1)

OpenStudy (dan815):

you have to solve for a1 first

OpenStudy (dan815):

27=a1*3^5 a1=27/3^5 =?

OpenStudy (dan815):

then sub a1 and r into the An formula

OpenStudy (anonymous):

so what is my n in the an formula?